OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form? Answer y = x – 4 y = x + 4 y = –x y = x

OpenStudy (anonymous):

your line will have slope 1 since it is perpendicular to the line with slope -1

OpenStudy (anonymous):

so the equation is \[y+2=1(x-2)\] \[y+2=x-2\] \[y=x-4\]